# Well-posedness of a first-order-in-time model for nonlinear acoustics with nonhomogeneous boundary conditions

### v2

@@ 1.1 Motivation
The motivation to study system (1) under the various boundary conditions (2a) - (2d) stems from the results established in [11]. The present work can be viewed as a continuation of the research direction initiated in [11], where the authors study well-posedness of the coupled nonlinear PDE
$\begin{split}\partial_{t}p+(1+\gamma)\nabla\cdot v-\zeta\Delta p+\epsilon_{1}p\nabla\cdot v+\epsilon_{2}\nabla p\cdot v&=p_{f}\\ \partial_{t}v+(1+\delta)\nabla p-\mu\Delta v-\frac{\epsilon_{3}}{2}\nabla(p^{2})+\frac{\epsilon_{4}}{2}\nabla(v\cdot v)&=v_{f}\\ p(\cdot,0)=p_{0},\quad v(\cdot,0)&=v_{0}\\ p=0,\quad v=0&\text{ on }\partial\Omega\end{split}$
with $\epsilon_{1},...,\epsilon_{4}>0$, and $\gamma,\delta\in L^{\infty}(\Omega)$ sufficiently small given, see [11, Eq. (13)]. Equation (1.1) can be interpreted as a reduced model derived from the Navier–Stokes–Fourier system. It serves as an approximation for nonlinear acoustic wave propagation in a viscous fluid, capturing the interaction between local fluctuations in pressure $p$ and the velocity field $v$, while not considering thermal and density variations. For a detailed derivation of (1.1) from fundamental physical principles, including the conservation of mass, momentum, and energy, as well as thermodynamic laws, we refer the reader to [11, Section 2].
@@
Moreover, as can be seen from [11, Remark 2.4], setting $\gamma=\delta=0$ in equation (1.1) yields a model that can be viewed as an alternative first-order-in-time formulation of Kuznetsov’s equation
$\begin{split}\partial_{tt}p-c^{2}\Delta p-b\partial_{t}\Delta p+\varepsilon\frac{B}{2A}\partial_{tt}(p^{2})-\varepsilon\partial_{tt}(v\cdot v)=p_{f}\\ \partial_{t}v=-\nabla p\end{split}$
originally derived in [15]. Here $c$ denotes the wave speed, $b$ is a damping parameter, $\varepsilon$ a small parameter and $\frac{B}{A}$ is the parameter of nonlinearity. Kuznetsov’s equation is widely used in nonlinear acoustics as a canonical model for high-intensity sound propagation. For a physical introduction to nonlinear acoustics, we refer to [7], and for a mathematical overview of the topic, we refer to the review article [8]. Unlike the model presented here, Kuznetsov’s equation includes second-order time derivatives, even within the nonlinear terms.
Interest in first-order formulations stems largely from their practical advantages in the design of efficient numerical methods, particularly in the context of adaptive space-time discretizations. In the case of the linear acoustics equation (1.1), effectiveness has already been demonstrated as a foundation for discontinuous Galerkin methods, see [11, Section 1] and the references therein.
@@
Before proceeding, we comment on the choice of the parameters used to derive (1) from the more general system (1.1). First, we set $\epsilon_{1}=\lambda,\epsilon_{2}=1,\epsilon_{3}=1,\epsilon_{3}=1,\epsilon_{4}=1$, a choice justified by the transform of variables $\tilde{p}=\varepsilon p$ and $\tilde{v}=\varepsilon v$, cf. [11, Eq. (12)]. In [11] the parameters $\epsilon_{1},\dots,\epsilon_{4}$ are treated as placeholders, allowing for flexibility in the coefficients of the nonlinear terms. However, in this work, we choose to simplify the model by avoiding additional parameters and instead focus on the core structure of the equation relevant to nonlinear acoustics. Second, we set $\gamma=\delta=0$, since, as noted earlier, Kuznetsov’s equation can be recovered from equation (1) under this assumption. Moreover, the terms involving $\gamma,\delta$ are related to entropy contributions, which are typically neglected in models of nonlinear acoustics. While the analysis presented here could, in principle, be extended to the case $\gamma,\delta\neq 0$ following similar arguments as in [11], we omit this extension to streamline the exposition and reduce notation.
@@ 1.2 Contributions
First, this paper extents the results of [11] by proving existence of $H^{1+s}(\Omega)$ regular solutions in space for $s\in[\frac{1}{2},1]$ in the case of nonhomogeneous Dirichlet boundary conditions, see Theorem 3.1. Compared to [11], the approach presented here accommodates slightly less regular initial data $(p_{0},v_{0})$ and right-hand sides $(p_{f},v_{f})$. An immediate consequence is that, for the limiting case $s=\frac{1}{2}$ solutions that are unbounded are not necessarily excluded from a theoretic standpoint, since $H^{s}(\Omega)\hookrightarrow L^{\infty}(\Omega)$ only for $s>\frac{3}{2}$ if $d=3$. To the author’s knowledge, fractional regularity results of this type have not been investigated for other equations in nonlinear acoustics.
@@
To obtain these results, we consider the more abstract formulations (2) and (2), which are semilinear parabolic evolution equations that have been extensively studied in the literature, see, for example, [16, 18].
For instance, well-posedness of (1) can be established using standard semigroup techniques, as presented in [18, 22]. By applying classical semigroup theory, well-posedness of equation (2) follows under the condition $g\in D(\mathcal{A}^{\gamma})$ with $\gamma>\frac{3}{4}$, see [22, Section 8.3, Theorem 3.5]. While there also exist results for less regular initial data, these typically require stronger assumptions on the right-hand side $f$, cf. [18, Theorem 7.1.6], while in this work the right hand side can be any function in $L^{2}$, thereby including discontinuities in both space and time.
Alternatively, well-posedness of equation (1) can be shown via energy methods, as in [16]. However, the analysis in [16] only considers fractional Hölder regularity and does not address fractional Sobolev regularity. While fractional Sobolev regularity results are provided in [24], those results are limited to the whole space $\mathbb{R}^{d}$. To the best knowledge of the author, a higher-order fractional Sobolev energy estimate on a bounded domain, combined with the specific type of nonlinearity considered here, has not yet been addressed in the literature.
@@ 1.3 Choice of boundary conditions
By Hodge/Lions boundary conditions we refer to the two boundary conditions involving the velocity $v$. In the Navier–Stokes literature, these are also known as Hodge, Lions, or Navier-slip boundary conditions, see, e.g., [17, 21, 23, 28]. Physically, such conditions are often used to model friction along the boundary of the domain. The specific Hodge/Lions formulation considered here is a special case of slip boundary conditions. This choice is particularly natural for equation (1), which is derived from the Navier–Stokes–Fourier system for viscous, heat-conducting fluids, cf. [11, Section 2].
If slip effects are not to be modeled, one may instead impose Dirichlet conditions on each component of $v$, leading to the classical no-slip condition $v=0$ on $\partial\Omega$. See [17] for an overview of different boundary conditions used for the Navier–Stokes equations. The present work therefore treats both slip and no-slip boundary conditions, while also allowing a choice between Dirichlet and Neumann conditions for $p$.
@@ 2 Abstract system
$ab\leq\frac{\delta a^{2}}{2}+\frac{b^{2}}{2\delta}\quad a,b,\delta>0$
@@
    $\langle\mathscr{A}u,u\rangle=0\quad\text{ for }u\in{D}(\mathfrak{A}).$
@@
    $|\langle\mathfrak{B}[u,w],v\rangle|\leq C_{\mathfrak{B}}\|\mathcal{A}^{\frac{\sigma+1}{2}}u\|_{H}\|\mathcal{A}^{\frac{\sigma}{2}}w\|_{H}\|\mathcal{A}^{\frac{1-\sigma}{2}}v\|_{H}$
@@
$\mathcal{A}^{s}u:{D}(\mathcal{A}^{s})\subset H\to H,\quad\mathcal{A}^{s}u:=\sum_{k=1}^{\infty}\lambda_{k}^{s}\langle u,\phi_{k}\rangle\phi_{k}$
@@
${D}(\mathcal{A}^{s})=\left\{u\in H:\sum_{k=1}^{\infty}\lambda_{k}^{2s}\langle u,\phi_{k}\rangle^{2}<\infty\right\},$
see [29, Section 5.8] for details.
@@
and therefore $\mathcal{A}^{s}:{D}(\mathcal{A}^{t})\to{D}(\mathcal{A}^{t-s})$ continuously for any $t\in\mathbb{R}$ cf. [6, Section 2.1].
@@
${D}(\mathcal{A}^{s})\hookrightarrow{D}(\mathcal{A}^{t})$
@@
${D}(\mathcal{A}^{\theta})=(H,{D}(\mathcal{A}))_{\theta}$
with $0<\theta<1$ denoting the parameter of real interpolation on Hilbert spaces, see [19, Theorem 4.36].
@@
$L^{2}(0,T;{D}(\mathcal{A}^{\frac{\sigma+1}{2}}))\cap H^{1}(0,T;{D}(\mathcal{A}^{\frac{1-\sigma}{2}})^{*})\hookrightarrow C([0,T];{D}(\mathcal{A}^{\frac{\sigma}{2}}))$
is continuous for any $T>0$ with embedding constant $C_{T}>0$ depending on $T$. This follows, since ${D}(\mathcal{A}^{\frac{\sigma}{2}})$ is a Hilbert space and from (2.1) we have ${D}(\mathcal{A}^{\frac{\sigma}{2}})\hookrightarrow{D}(\mathcal{A}^{\frac{\sigma+1}{2}})$, yielding the Gelfand tripel ${D}(\mathcal{A}^{\frac{\sigma}{2}})\subset{D}(\mathcal{A}^{\frac{\sigma+1}{2}})\subset{D}(\mathcal{A}^{\frac{\sigma+1}{2}})^{*}$. Applying [25, Lemma 7.3, $p=p^{\prime}=2$] yields $L^{2}(0,T;{D}(\mathcal{A}^{\frac{\sigma+1}{2}}))\cap H^{1}(0,T;{D}(\mathcal{A}^{\frac{1+\sigma}{2}})^{*})\hookrightarrow C([0,T];{D}(\mathcal{A}^{\frac{\sigma}{2}}))$. The embedding (2.1) then follows since, ${D}(\mathcal{A}^{\frac{\sigma-1}{2}})^{*}={D}(\mathcal{A}^{\frac{-\sigma+1}{2}})\hookrightarrow{D}(\mathcal{A}^{\frac{-\sigma-1}{2}})={D}(\mathcal{A}^{\frac{\sigma+1}{2}})^{*}$.
@@ 2.1 Auxiliary linear problem
$\|u\|_{X}\leq C_{G}(\|f\|_{Y}+\|g\|_{{D}(\mathcal{A}^{\frac{\sigma}{2}})})$
@@
First, we show that $\mathfrak{A}:X\to Y$ is bounded. Due to assumption 2.1(ii) and embedding (2.1), we know that $\mathcal{A}:U\to V^{*}$ is bounded with some constant $C_{\mathcal{A}}>0$, since $U$ and $V^{*}$ can be expressed in terms of fractional domains of $\mathcal{A}$. Furthermore, by boundedness of $\mathscr{A}:{D}(\mathcal{A}^{\frac{1}{2}})\to H$ and (2.1) from Remark 2.1
@@
Second, we have boundedness of $\mathfrak{B}:X\times X\to Y$ with constant less or equal then $C_{\mathfrak{B}}C_{T}$. This is due to assumption 2.1(iii) implying (2.1(iii)) and (2.1), since then
$\int_{0}^{T}\|\mathfrak{B}[u,v]\|^{2}_{V^{*}}ds\leq\int_{0}^{T}C_{\mathfrak{B}}^{2}\|u\|_{U}^{2}\|v\|^{2}_{W}ds\leq\max_{s\in[0,T]}\|v\|_{W}^{2}C_{\mathfrak{B}}^{2}\int_{0}^{T}\|u\|_{U}^{2}ds\leq C_{T}^{2}C_{\mathfrak{B}}^{2}\|u\|_{X}^{2}\|v\|_{X}^{2}.$

### v3

@@ 1.1 Motivation
The motivation to study system (1) under the various boundary conditions (2a) - (2d) stems from the results established in [QuadraticWave]. The present work can be viewed as a continuation of the research direction initiated in [QuadraticWave], where the authors study well-posedness of the coupled nonlinear PDE
$\begin{split}\partial_{t}p+(1+\gamma)\nabla\cdot v-\zeta\Delta p+\epsilon_{1}p\nabla\cdot v+\epsilon_{2}\nabla p\cdot v&=p_{f}\\ \partial_{t}v+(1+\delta)\nabla p-\mu\Delta v-\frac{\epsilon_{3}}{2}\nabla(p^{2})+\frac{\epsilon_{4}}{2}\nabla(v\cdot v)&=v_{f}\\ p(\cdot,0)=p_{0},\quad v(\cdot,0)&=v_{0}\\ p=0,\quad v&=0\text{ on }\partial\Omega\end{split}$
with $\epsilon_{1},...,\epsilon_{4}>0$, and $\gamma,\delta\in L^{\infty}(\Omega)$ sufficiently small given, see [QuadraticWave, Eq. (13)]. Equation (1.1) can be interpreted as a reduced model derived from the Navier–Stokes–Fourier system. It serves as an approximation for nonlinear acoustic wave propagation in a viscous fluid, capturing the interaction between local fluctuations in pressure $p$ and the velocity field $v$, while not considering thermal and density variations. For a detailed derivation of (1.1) from fundamental physical principles, including the conservation of mass, momentum, and energy, as well as thermodynamic laws, we refer the reader to [QuadraticWave, Section 2].
@@
Moreover, as observed in [QuadraticWave, Remark 2.4], setting $\gamma=\delta=0$ in equation (1.1) yields a model that can be viewed as an alternative first-order-in-time formulation of Kuznetsov’s equation. The latter is given by
$\begin{split}\partial_{tt}p-c^{2}\Delta p-b\partial_{t}\Delta p-\varepsilon\frac{B}{2A}\partial_{tt}(p^{2})-\varepsilon\partial_{tt}(v\cdot v)&=0\\ \partial_{t}v+\textasciiacute\nabla p&=0\\ p(\cdot,0)=p_{0},\quad\partial_{t}p(\cdot,0)=p_{1},\quad v(\cdot,0)&=v_{0}\\ p&=0\text{ on }\partial\Omega\end{split}$
and originally derived in [kuznetsov:1971]. Here $c$ denotes the wave speed, $b$ is a damping parameter, $\varepsilon$ a small parameter and $\frac{B}{A}$ is the parameter of nonlinearity. Kuznetsov’s equation is widely used in nonlinear acoustics as a canonical model for high-intensity sound propagation. For a physical introduction to nonlinear acoustics, we refer to [Hamilton:1997], while a comprehensive mathematical overview can be found in the review article [Kaltenbacher:2015].
In contrast to the model (1) considered in the present work, Kuznetsov’s equation involves second-order time derivatives, including within the nonlinear terms. Consequently, the regularity theory differs from that developed here, see, [Kaltenbacher:2015, MeyerWilke2013]. In particular, [MeyerWilke2013, Theorem 1] establishes optimal $L^{p}$-regularity results for (1.1). In the Hilbert space setting $p=2$, assuming $p_{0}\in H^{2}(\Omega)$, $p_{1},v_{0}\in H^{1}(\Omega)$, the authors obtain $p\in H^{2}(0,T;L^{2}(\Omega))\cap H^{1}(0,T;H^{2}(\Omega))$, $v\in BUC^{1}(0,T;H^{1}(\Omega))$ and $\partial_{t}v\in H^{\frac{3}{2}}(0,T;L^{2}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega))$. Here, $BUC^{1}$ denotes the space of functions having a bounded and uniformly continuous derivative.
Interest in first-order formulations stems largely from their practical advantages in the design of efficient numerical methods, particularly in the context of adaptive space-time discretizations. In the case of the linear acoustics equation (1.1), effectiveness has already been demonstrated as a foundation for discontinuous Galerkin methods, see [DoerflerFindeisenWienersZiegler:2019, bansal:2021].
@@
Before proceeding, we comment on the choice of the parameters used to derive (1) from the more general system (1.1). First, we set $\epsilon_{1}=\lambda,\epsilon_{2}=1,\epsilon_{3}=1,\epsilon_{3}=1,\epsilon_{4}=1$, a choice justified by the transform of variables $\tilde{p}=\varepsilon p$ and $\tilde{v}=\varepsilon v$, cf. [QuadraticWave, Eq. (12)]. In [QuadraticWave] the parameters $\epsilon_{1},\dots,\epsilon_{4}$ are treated as placeholders, allowing for flexibility in the coefficients of the nonlinear terms. However, in this work, we choose to simplify the model by avoiding additional parameters and instead focus on the core structure of the equation relevant to nonlinear acoustics. Second, we set $\gamma=\delta=0$, since, as noted earlier, Kuznetsov’s equation can be recovered from equation (1) under this assumption. Moreover, the terms involving $\gamma,\delta$ are related to entropy contributions, which are typically neglected in models of nonlinear acoustics. While the analysis presented here could, in principle, be extended to the case $\gamma,\delta\neq 0$ following similar arguments as in [QuadraticWave], we omit this extension to streamline the exposition and reduce notation.
@@ 1.2 Contributions
First, this paper extends the results of [QuadraticWave] by proving existence of $H^{1+s}(\Omega)$ regular solutions in space for $s\in[\frac{1}{2},1]$ in the case of nonhomogeneous Dirichlet boundary conditions, see Theorem 3.1. Compared to [QuadraticWave], the approach presented here accommodates slightly less regular initial data $(p_{0},v_{0})$ and right-hand sides $(p_{f},v_{f})$. An immediate consequence is that, for the limiting case $s=\frac{1}{2}$ solutions that are unbounded are not necessarily excluded from a theoretic standpoint, since $H^{s}(\Omega)\hookrightarrow L^{\infty}(\Omega)$ only for $s>\frac{3}{2}$ if $d=3$. To the author’s knowledge, fractional regularity results of this type have not been investigated for other equations in nonlinear acoustics.
@@
To obtain these results, we consider the more abstract formulations (2) and (2), which are semilinear parabolic evolution equations that have been extensively studied in the literature, see, for example, [Ladyzhenskaya1968, lunardi1995].
For instance, well-posedness of (1) can be established using standard semigroup techniques, as presented in [lunardi1995, pazy1983]. By applying classical semigroup theory, well-posedness of equation (2) follows under the condition $g\in D(\mathcal{A}^{\gamma})$ with $\gamma>\frac{3}{4}$, see [pazy1983, Section 8.3, Theorem 3.5]. While there also exist results for less regular initial data, these typically require stronger assumptions on the right-hand side $f$, cf. [lunardi1995, Theorem 7.1.6], while in this work the right hand side can be any function in $L^{2}$, thereby including discontinuities in both space and time.
Alternatively, well-posedness of equation (1) can be shown via energy methods, as in [Ladyzhenskaya1968]. However, the analysis in [Ladyzhenskaya1968] only considers fractional Hölder regularity and does not address fractional Sobolev regularity. While fractional Sobolev regularity results are provided in [Ribaud1998], those results are limited to the whole space $\mathbb{R}^{d}$. To the best knowledge of the author, a higher-order fractional Sobolev energy estimate on a bounded domain, combined with the specific type of nonlinearity considered here, has not yet been addressed in the literature.
@@ 1.3 Choice of boundary conditions
By Hodge/Lions boundary conditions we refer to the two boundary conditions involving the velocity $v$. In the Navier–Stokes literature, these are also known as Hodge, Lions, or Navier-slip boundary conditions, see, e.g., [li2022b, monniaux2013, phan2017, xiao2007]. Physically, such conditions are often used to model friction along the boundary of the domain. The specific Hodge/Lions formulation considered here is a special case of slip boundary conditions. This choice is particularly natural for equation (1), which is derived from the Navier–Stokes–Fourier system for viscous, heat-conducting fluids, cf. [QuadraticWave, Section 2].
If slip effects are not to be modeled, one may instead impose Dirichlet conditions on each component of $v$, leading to the classical no-slip condition $v=0$ on $\partial\Omega$. See [li2022b] for an overview of different boundary conditions used for the Navier–Stokes equations. The present work therefore treats both slip and no-slip boundary conditions, while also allowing a choice between Dirichlet and Neumann conditions for $p$.
@@ 2 Abstract system
$ab\leq\frac{\delta a^{2}}{2}+\frac{b^{2}}{2\delta}\quad a,b,\delta>0$ (9)
@@
    $\langle\mathscr{A}u,u\rangle=0\quad\text{ for }u\in{D}(\mathfrak{A}).$ (10)
@@
    $|\langle\mathfrak{B}[u,w],v\rangle|\leq C_{\mathfrak{B}}\|\mathcal{A}^{\frac{\sigma+1}{2}}u\|_{H}\|\mathcal{A}^{\frac{\sigma}{2}}w\|_{H}\|\mathcal{A}^{\frac{1-\sigma}{2}}v\|_{H}$ (11)
@@
$\mathcal{A}^{s}u:{D}(\mathcal{A}^{s})\subset H\to H,\quad\mathcal{A}^{s}u:=\sum_{k=1}^{\infty}\lambda_{k}^{s}\langle u,\phi_{k}\rangle\phi_{k}$ (12)
@@
${D}(\mathcal{A}^{s})=\left\{u\in H:\sum_{k=1}^{\infty}\lambda_{k}^{2s}\langle u,\phi_{k}\rangle^{2}<\infty\right\},$ (13)
see [zeidler1995, Section 5.8] for details.
@@
and therefore $\mathcal{A}^{s}:{D}(\mathcal{A}^{t})\to{D}(\mathcal{A}^{t-s})$ continuously for any $t\in\mathbb{R}$ cf. [fefferman2022, Section 2.1].
@@
${D}(\mathcal{A}^{s})\hookrightarrow{D}(\mathcal{A}^{t})$ (14)
@@
${D}(\mathcal{A}^{\theta})=(H,{D}(\mathcal{A}))_{\theta}$ (15)
with $0<\theta<1$ denoting the parameter of real interpolation on Hilbert spaces, see [Lunardi2018, Theorem 4.36].
@@
$L^{2}(0,T;{D}(\mathcal{A}^{\frac{\sigma+1}{2}}))\cap H^{1}(0,T;{D}(\mathcal{A}^{\frac{1-\sigma}{2}})^{*})\hookrightarrow C([0,T];{D}(\mathcal{A}^{\frac{\sigma}{2}}))$ (16)
is continuous for any $T>0$ with embedding constant $C_{T}>0$ depending on $T$. This follows, since ${D}(\mathcal{A}^{\frac{\sigma}{2}})$ is a Hilbert space and from (14) we have ${D}(\mathcal{A}^{\frac{\sigma}{2}})\hookrightarrow{D}(\mathcal{A}^{\frac{\sigma+1}{2}})$, yielding the Gelfand tripel ${D}(\mathcal{A}^{\frac{\sigma}{2}})\subset{D}(\mathcal{A}^{\frac{\sigma+1}{2}})\subset{D}(\mathcal{A}^{\frac{\sigma+1}{2}})^{*}$. Applying [Roubicek:2013, Lemma 7.3, $p=p^{\prime}=2$] yields $L^{2}(0,T;{D}(\mathcal{A}^{\frac{\sigma+1}{2}}))\cap H^{1}(0,T;{D}(\mathcal{A}^{\frac{1+\sigma}{2}})^{*})\hookrightarrow C([0,T];{D}(\mathcal{A}^{\frac{\sigma}{2}}))$. The embedding (16) then follows since, ${D}(\mathcal{A}^{\frac{\sigma-1}{2}})^{*}={D}(\mathcal{A}^{\frac{-\sigma+1}{2}})\hookrightarrow{D}(\mathcal{A}^{\frac{-\sigma-1}{2}})={D}(\mathcal{A}^{\frac{\sigma+1}{2}})^{*}$.
@@ 2.1 Auxiliary linear problem
$\|u\|_{X}\leq C_{G}(\|f\|_{Y}+\|g\|_{{D}(\mathcal{A}^{\frac{\sigma}{2}})})$ (17)
@@
First, we show that $\mathfrak{A}:X\to Y$ is bounded. Due to assumption 2.1(ii) and embedding (14), we know that $\mathcal{A}:U\to V^{*}$ is bounded with some constant $C_{\mathcal{A}}>0$, since $U$ and $V^{*}$ can be expressed in terms of fractional domains of $\mathcal{A}$. Furthermore, by boundedness of $\mathscr{A}:{D}(\mathcal{A}^{\frac{1}{2}})\to H$ and (14) from Remark 2.1
@@
Second, we have boundedness of $\mathfrak{B}:X\times X\to Y$ with constant less or equal then $C_{\mathfrak{B}}C_{T}$. This is due to assumption 2.1(iii) implying (11) and (16), since then
$\int_{0}^{T}\|\mathfrak{B}[u,v]\|^{2}_{V^{*}}ds\leq\int_{0}^{T}C_{\mathfrak{B}}^{2}\|u\|_{U}^{2}\|v\|^{2}_{W}ds\leq\max_{s\in[0,T]}\|v\|_{W}^{2}C_{\mathfrak{B}}^{2}\int_{0}^{T}\|u\|_{U}^{2}ds\leq C_{T}^{2}C_{\mathfrak{B}}^{2}\|u\|_{X}^{2}\|v\|_{X}^{2}.$ (18)